# Non-Asymptotic Analysis of Efficiency in Conformalized Regression

### v1

@@ 1 Introduction
Deploying machine learning models in safety-critical domains, such as health care (Allgaier et al., 2023; Gui et al., 2024), finance (Wisniewski et al., 2020; Bastos, 2024), and autonomous systems (Lindemann et al., 2023; Ren et al., 2023), requires not only accurate predictions but also reliable uncertainty quantification. Conformal prediction (CP) is a principled, distribution-free framework for this purpose, equipping black-box models with prediction sets achieving coverage guarantees or validity (Vovk et al., 2005; Balasubramanian et al., 2014). Formally, given a set of data $\{(X_{j},Y_{j})\}_{j=1}^{m}$ drawn from a distribution $\mathcal{P}$ over $\mathcal{X}\times\mathcal{Y}$, for any user-specified miscoverage level $\alpha\in(0,1)$ and a predictive model, conformal prediction constructs a set-valued function $\mathcal{C}:\mathcal{X}\to 2^{\mathcal{Y}}$ such that, for a test pair $(X_{m+1},Y_{m+1})\sim\mathcal{P}$, the prediction set $\mathcal{C}(X_{m+1})$ covers the label $Y_{m+1}$ with probability
@@
A central focus of conformal prediction is efficiency, commonly quantified by the expected measure of the prediction set (Shafer and Vovk, 2008). For classification tasks, efficiency relates to the cardinality of the predicted label set; for regression, it corresponds to the length (or volume) of the prediction interval (or region). Under the validity condition (1), smaller prediction sets are more informative. Early works primarily evaluated efficiency empirically, whereas recent research has shifted toward asymptotic efficiency, demonstrating that prediction sets converge to the oracle sets as the sample size increases (Sesia and Candès, 2020; Chernozhukov et al., 2021; Izbicki et al., 2022). In contrast, non-asymptotic efficiency, or finite-sample guarantees on the expected measure or excess measure of the prediction set, remains much less understood, with only partial results available (Lei and Wasserman, 2014; Lei et al., 2018; Dhillon et al., 2024; Bars and Humbert, 2025). Existing non-asymptotic bounds are typically expressed based on the calibration set size $m$, whereas the effect of training set size $n$ and miscoverage level $\alpha$ remains an open question in split conformalized regression.
In this work, we analyze the efficiency of split conformal prediction in regression, focusing on conformalized median regression (CMR) and conformalized quantile regression (CQR) (Romano et al., 2019). CMR uses the absolute residual as the nonconformity score, and the quantile of the calibration residuals then determines the half-width of a symmetric prediction interval centered at the estimated conditional median. In contrast, CQR estimates both upper and lower conditional quantiles, defining nonconformity scores relative to these estimates. After calibration, CQR yields adaptive, asymmetric prediction intervals that naturally capture heteroscedasticity without assuming symmetric conditional quantiles.
@@ 3.1 Problem Setup for CQR-SGD
In CQR, the training set $\mathcal{D}_{\mathrm{train}}$ is used to estimate the conditional $\gamma$-quantile function $q_{\gamma}\left(Y\mid X\right)$ defined in (2), where $\gamma=1-\alpha/2,\alpha/2$. The estimated function $t_{\gamma}(\cdot;\theta_{n}(\gamma))$ is obtained by solving the stochastic pinball loss minimization problem (Koenker and Bassett Jr, 1978):
@@
Without loss of generality, we assume $K\leq\max\{|y_{\min}|,|y_{\max}|\}/B$. The linear model represents a standard setting for theoretical analysis of quantile regression (Koenker, 2005; Pan and Zhou, 2021), ensuring convexity of the objective function in (4).
@@
To solve (4), we consider the stochastic approximation framework (Robbins and Monro, 1951), focusing on stochastic gradient descent (SGD). The $\theta$ is updated according to
@@
The phenomenon where the lower quantile estimate exceeds the upper quantile estimate is known as quantile crossing (Romano et al., 2019; Bassett Jr and Koenker, 1982). We show in the proof of Proposition B.8 that, quantile crossing does not occur with high probability once the training set size $n$ is sufficiently large.
@@ 3.2 Theoretical Results for Efficiency of CQR
$q_{\gamma}(Y\mid X=x)=t_{\gamma}(x;\theta^{*}(\gamma))\quad\text{for all }x\in\mathcal{X}.$
Assumption 3.1 ensures that $\theta^{*}(\gamma)$ is a minimizer of (4) (Takeuchi et al., 2006; Steinwart and Christmann, 2011).
@@
If Assumptions 3.13.3 hold, then
@@
The proof of Theorem 3.1 is deferred to Appendix B.1.
@@
The results of Theorem 3.1 are established under a strongly-convex assumption as they rely on Theorem B.4 from Rakhlin et al. (2012). Comparable rates can also be obtained for non-strongly-convex objectives under the assumptions in Bach and Moulines (2013), where Assumption 3.2 can be weakened to requiring only the invertibility of $\mathbb{E}[XX^{\top}]$.
Remark 3.3.
Faster rates than those of Theorem 3.1 are attainable with variance-reduced stochastic gradients; see Appendix A for further discussion.
@@
The explicit upper bound (B.5) and the full proof of Theorem 3.2 are presented in Appendix C, with a proof sketch illustrated in Figure 1.
@@
Figure 1: Proof outline of Theorem 3.2. Full proof deferred to Section B.
@@
For a random variable $Z$, the density level set $\mathcal{L}(u_{1-\alpha})$ is the optimal prediction set with coverage probability $1-\alpha$ (Lei et al., 2011), i.e.,
@@ 3.2.1 Phase Transitions of the Upper Bound
  • Thus, if $\alpha=$ $\Omega(\max\{n^{-1/4},\sqrt{\log m/m}\})$, the upper bound scales as $\mathcal{O}(n^{-1/2}+m^{-1/2})$, which coincides with the rate in Bars and Humbert (2025) assuming a finite function class.
@@
Figure 2: Upper bound orders in Theorem 3.2 in different regimes of $\alpha$ when $n=\Theta(m)$. Results in Lei et al. (2018); Bars and Humbert (2025) lie in the right most regime (blue).
@@ 4.2 Theoretical Results for Efficiency of CMR
Assumption 4.2 is standard in the analysis of conformalized regression based on a single regressor, following the precedent set by Assumption A1 of Lei et al. (2018).
@@
The explicit upper bound (C) and the full proof of Theorem 4.1 are presented in Appendix C.
@@ 5 Related Works
Quantile regression has attracted significant attention since the seminal work of Koenker and Bassett Jr (1978) due to its robustness to outliers and ability to capture distributional heterogeneity. Early works derived the $\sqrt{n}$-consistency and asymptotic normality of quantile regressors in the linear model (Bassett Jr and Koenker, 1978, 1982; Portnoy and Koenker, 1989; Pollard, 1991). Other works established statistical properties under fixed designs, where covariates are treated as deterministic (He and Shao, 1996; Koenker, 2005). More recent works have shifted toward non-asymptotic analysis with convergence rate $\mathcal{O}(1/\sqrt{n})$ under random designs, where covariates are random and prediction performance on unseen data is emphasized (Steinwart and Christmann, 2011; Catoni, 2012; Hsu et al., 2014; Loh and Wainwright, 2013; Pan and Zhou, 2021; He et al., 2023; Liu et al., 2023; Sasai and Fujisawa, 2025). Median regression is a special case of quantile regression, has also been extensively studied (Chen et al., 2008). These methods form the basis for conformalized median regression and conformalized quantile regression (Romano et al., 2019).
@@
Conformal prediction was developed to equip point predictions with confidence regions that provide finite-sample coverage guarantees (Papadopoulos et al., 2002; Vovk et al., 2005, 2009; Vovk, 2025). Research on its efficiency (Vovk et al., 2016; Gasparin and Ramdas, 2025) has evolved from early asymptotic convergence analyses, which established convergence rates toward the oracle prediction region (Chajewska et al., 2001; Li and Liu, 2008; Sadinle et al., 2019; Sesia and Candès, 2020; Chernozhukov et al., 2021; Izbicki et al., 2022), to generalization error-based bounds on expected set size Zecchin et al. (2024), and recently volume-minimization methods using data-driven norms (Sharma et al., 2023; Correia et al., 2024; Kiyani et al., 2024; Braun et al., 2025; Bars and Humbert, 2025; Gao et al., 2025; Srinivas, 2025).
For conditional density estimation, under $\beta$-Hölder class and $\gamma$-exponent margin conditions of the conditional density, Lei and Wasserman (2014) derived minimax-optimal rates of order $\mathcal{O}((\log m/m)^{\beta/(3\beta+1)})$ when $\gamma=1$, and showed that conditional coverage cannot generally be guaranteed in finite samples. When the quantile of $Y$ is symmetric and independent of $X$ (analogous to Assumption 4.2), Lei et al. (2018) incorporated training error into the efficiency analysis, treating $\alpha$ as a fixed constant. In contrast, our results for CQR and CMR make no assumptions on the training error and provide explicit upper bounds (B.5, C) as functions of $(n,m,\alpha)$, applicable also to adaptive prediction sets.
Under the assumptions that the quantile function of the nonconformity score is locally $\beta$-Hölder continuous, and that the worst-case empirical estimation error of the function class is bounded, Bars and Humbert (2025) derived convergence rates of the order $\mathcal{O}(m^{-\beta\kappa/2}+n^{-\beta\iota/2})$ for some $0<\iota,\kappa<1$ when the function class is finite. In the case of $\beta=1$, this rate matches our bound when $\alpha$ is treated as a fixed constant, namely $\mathcal{O}(m^{-1/2}+n^{-1/2})$. Different from analysis in Bars and Humbert (2025) that focuses on methods based on volume minimization, our work develops efficiency guarantees for CQR and CMR, without imposing assumptions on the score distribution induced by the trained model or on the estimation error. Instead, we demonstrate in the proof (especially Proposition C.2) that the required regularity conditions of the score are satisfied with high probability under mild assumptions on the underlying data distribution.
@@ 6 Experiments
This section presents evaluations of length deviation using synthetic data to access our theoretical results. Real-world experiments are deferred to Appendix E due to space constraints.
Experiment setup. The data generation procedure is described in Appendix D.1. All experiments employ linear models trained with SGD for one epoch using a batch size of $64$. Learning rates are selected via successive halving over the range $10^{-5},1]$. We evaluate miscoverage levels $\alpha\in\{0.01,0.025,0.05,0.075,0.1,0.125,0.15,0.175,0.2\}$. Reported results are averaged over $20$ independent trials, and length deviations are computed on $2000$ test samples.
@@
    Together with (23), the estimated coefficient $b_{1}=-2.24$ (Fig. 3d) implies that $\Delta\sim\alpha^{-2.24}$. This aligns with the theoretical upper bound of order $\mathcal{O}(\alpha^{-2})$. Appendix D.2 provides an additional verification for the existence of this regime.
@@ 7 Conclusion
Appendix A Discussion on Upper bounds for Quantile Estimation Error
In Theorem 3.1, we provide the upper bound $\mathcal{O}(n^{-1})$ for the quantile estimation error using standard SGD when the objective (4) is strongly convex under Assumption 3.2 and 3.3. We notice that Assumption 3.2 can be relaxed as discussed in Remark 3.2.
We also note that variance-reduction techniques, such as SAG (Schmidt et al., 2017), SVRG (Johnson and Zhang, 2013), and SAGA (Defazio et al., 2014), achieve a linear convergence rate under strong convexity, meaning that the error decays geometrically in the number of iterations $\mathcal{O}(\exp(-cn))$. This is in sharp contrast to the $\mathcal{O}(n^{-1})$ rate of standard SGD. Since our focus in this paper is not on improving the convergence rate of quantile estimation, we present results under the standard SGD rate in Theorem 3.1. Nevertheless, accelerated (linear) rates can be obtained by incorporating variance-reduction techniques following the same proof strategy developed here.
Appendix B Proofs of Results in CQR

### v2

@@ 1 Introduction
Deploying machine learning models in safety-critical domains, such as health care (allgaier2023does; gui2024conformal), finance (wisniewski2020application; bastos2024conformal), and autonomous systems (lindemann2023safe; ren2023robots), requires not only accurate predictions but also reliable uncertainty quantification. Conformal prediction (CP) is a principled, distribution-free framework for this purpose, equipping black-box models with prediction sets achieving coverage guarantees or validity (vovk2005algorithmic; balasubramanian2014conformal). Formally, given a set of data $\{(X_{j},Y_{j})\}_{j=1}^{m}$ drawn from a distribution $\mathcal{P}$ over $\mathcal{X}\times\mathcal{Y}$, for any user-specified miscoverage level $\alpha\in(0,1)$ and a predictive model, conformal prediction constructs a set-valued function $\mathcal{C}:\mathcal{X}\to 2^{\mathcal{Y}}$ such that, for a test pair $(X_{m+1},Y_{m+1})\sim\mathcal{P}$, the prediction set $\mathcal{C}(X_{m+1})$ covers the label $Y_{m+1}$ with probability
@@
A central focus of conformal prediction is efficiency, commonly quantified by the expected measure of the prediction set (shafer2008tutorial). For classification tasks, efficiency relates to the cardinality of the predicted label set; for regression, it corresponds to the length (or volume) of the prediction interval (or region). Under the validity condition (1), smaller prediction sets are more informative. Early works primarily evaluated efficiency empirically, whereas recent research has shifted toward asymptotic efficiency, demonstrating that prediction sets converge to the oracle sets as the sample size increases (sesia2020comparison; chernozhukov2021distributional; izbicki2022cd). In contrast, non-asymptotic efficiency, or finite-sample guarantees on the expected measure or excess measure of the prediction set, remains much less understood, with only partial results available (lei2014distribution; lei2018distribution; dhillon2024expected; bars2025volume). Existing non-asymptotic bounds are typically expressed based on the calibration set size $m$, whereas the effect of training set size $n$ and miscoverage level $\alpha$ remains an open question in split conformalized regression.
In this work, we analyze the efficiency of split conformal prediction in regression, focusing on conformalized median regression (CMR) and conformalized quantile regression (CQR) (romano2019conformalized). CMR uses the absolute residual as the nonconformity score, and the quantile of the calibration residuals then determines the half-width of a symmetric prediction interval centered at the estimated conditional median. In contrast, CQR estimates both upper and lower conditional quantiles, defining nonconformity scores relative to these estimates. After calibration, CQR yields adaptive, asymmetric prediction intervals that naturally capture heteroscedasticity without assuming symmetric conditional quantiles.
@@ 3.1 Problem Setup for CQR-SGD
In CQR, the training set $\mathcal{D}_{\mathrm{train}}$ is used to estimate the conditional $\gamma$-quantile function $q_{\gamma}\left(Y\mid X\right)$ defined in (2), where $\gamma=1-\alpha/2,\alpha/2$. The estimated function $t_{\gamma}(\cdot;\theta_{n}(\gamma))$ is obtained by solving the stochastic pinball loss minimization problem (koenker1978regression):
@@
Without loss of generality, we assume $K\leq\max\{|y_{\min}|,|y_{\max}|\}/B$. The linear model represents a standard setting for theoretical analysis of quantile regression (koenker2005quantile; pan2021multiplier), ensuring convexity of the objective function in (4).
@@
To solve (4), we consider the stochastic approximation framework (robbins1951stochastic), focusing on stochastic gradient descent (SGD). The $\theta$ is updated according to
@@
The phenomenon where the lower quantile estimate exceeds the upper quantile estimate is known as quantile crossing (romano2019conformalized; bassett1982empirical). We show in the proof of Proposition A.8 that, quantile crossing does not occur with high probability once the training set size $n$ is sufficiently large. Moreover, because the covariate space $\mathcal{X}$ is bounded, the ground-truth lower and upper quantile functions cannot cross, even if they are not parallel.
@@ 3.2 Theoretical Results for Efficiency of CQR
$q_{\gamma}(Y\mid X=x)=t_{\gamma}(x;\theta^{*}(\gamma))=x^{\top}\theta^{*}(\gamma),\quad\text{for all }x\in\mathcal{X}\ \subset\mathbb{R}^{d}.$
Assumption 3.1 ensures that $\theta^{*}(\gamma)$ is a minimizer of (4) (takeuchi2006nonparametric; steinwart2011estimating).
@@
If Assumptions 3.13.3 hold, taking step size $\eta_{k}=1/\left(\lambda_{\min}f_{\min}k\right)$ in SGD update (7), then
@@
The proof of Theorem 3.1 is deferred to Appendix A.1.
@@
The results of Theorem 3.1 are established under a strongly-convex assumption as they rely on Theorem A.4 from rakhlin2012making. Comparable rates can also be obtained for non-strongly-convex objectives under the assumptions in bach2013non, where Assumption 3.2 can be weakened to requiring only the invertibility of $\mathbb{E}[XX^{\top}]$.
@@
The explicit upper bound (A.5) and the full proof of Theorem 3.2 are presented in Appendix B, with a proof sketch illustrated in Figure 1.
@@
Figure 1: Proof outline of Theorem 3.2. Full proof deferred to Section A.
@@
For a random variable $Z$, the density level set $\mathcal{L}(u_{1-\alpha})$ is the optimal prediction set with coverage probability $1-\alpha$ (lei2011efficient), i.e.,
@@ 3.2.1 Phase Transitions of the Upper Bound
  • Thus, if $\alpha=$ $\Omega(\max\{n^{-1/4},\sqrt{\log m/m}\})$, the upper bound scales as $\mathcal{O}(n^{-1/2}+m^{-1/2})$, which coincides with the rate in bars2025volume assuming a finite function class.
@@
Figure 2: Upper bound orders in Theorem 3.2 in different regimes of $\alpha$ when $n=\Theta(m)$. Results in lei2018distribution; bars2025volume lie in the right most regime (blue).
@@ 4.2 Theoretical Results for Efficiency of CMR
Assumption 4.2 is standard in the analysis of conformalized regression based on a single regressor, following the precedent set by Assumption A1 of lei2018distribution.
@@
The explicit upper bound (B) and the full proof of Theorem 4.1 are presented in Appendix B.
@@ 5 Related Works
Quantile regression has attracted significant attention since the seminal work of koenker1978regression due to its robustness to outliers and ability to capture distributional heterogeneity. Early works derived the $\sqrt{n}$-consistency and asymptotic normality of quantile regressors in the linear model (bassett1978asymptotic; bassett1982empirical; portnoy1989adaptive; pollard1991asymptotics). Other works established statistical properties under fixed designs, where covariates are treated as deterministic (he1996general; koenker2005quantile). More recent works have shifted toward non-asymptotic analysis with convergence rate $\mathcal{O}(1/\sqrt{n})$ under random designs, where covariates are random and prediction performance on unseen data is emphasized (steinwart2011estimating; catoni2012challenging; hsu2014random; loh2015regularized; pan2021multiplier; he2023smoothed; liu2023distribution; sasai2025outlier). Median regression is a special case of quantile regression, has also been extensively studied (chen2008analysis). shen2024online analyze online quantile regression with linear models trained via SGD, under regularity conditions closely related to ours, including a local lower bound on the conditional density. These methods form the basis for conformalized median regression and conformalized quantile regression (romano2019conformalized).
@@
Conformal prediction was developed to equip point predictions with confidence regions that provide finite-sample coverage guarantees (papadopoulos2002inductive; vovk2005algorithmic; vovk2009line; vovk2025randomness). Research on its efficiency (vovk2016criteria; gasparin2025improving) has evolved from early asymptotic convergence analyses, which established convergence rates toward the oracle prediction region (chajewska2001learning; li2008multivariate; sadinle2019least; sesia2020comparison; chernozhukov2021distributional; izbicki2022cd), to generalization error-based bounds on expected set size zecchin2024generalization, and recently volume-minimization methods using data-driven norms (sharma2023pac; correia2024information; kiyani2024length; braun2025minimum; bars2025volume; gao2025volume; srinivas2025online).
For conditional density estimation, under $\beta$-Hölder class and $\gamma$-exponent margin conditions of the conditional density, lei2014distribution derived minimax-optimal rates of order $\mathcal{O}((\log m/m)^{\beta/(3\beta+1)})$ when $\gamma=1$, and showed that conditional coverage cannot generally be guaranteed in finite samples. When the quantile of $Y$ is symmetric and independent of $X$ (analogous to Assumption 4.2), lei2018distribution incorporated training error into the efficiency analysis, treating $\alpha$ as a fixed constant. In contrast, our results for CQR and CMR make no assumptions on the training error and provide explicit upper bounds (A.5, B) as functions of $(n,m,\alpha)$, applicable also to adaptive prediction sets.
Under the assumptions that the quantile function of the nonconformity score is locally $\beta$-Hölder continuous, and that the worst-case empirical estimation error of the function class is bounded, bars2025volume derived convergence rates of the order $\mathcal{O}(m^{-\beta\kappa/2}+n^{-\beta\iota/2})$ for some $0<\iota,\kappa<1$ when the function class is finite. In the case of $\beta=1$, this rate matches our bound when $\alpha$ is treated as a fixed constant, namely $\mathcal{O}(m^{-1/2}+n^{-1/2})$. Different from analysis in bars2025volume that focuses on methods based on volume minimization, our work develops efficiency guarantees for CQR and CMR, without imposing assumptions on the score distribution induced by the trained model or on the estimation error. Instead, we demonstrate in the proof (especially Proposition B.2) that the required regularity conditions of the score are satisfied with high probability under mild assumptions on the underlying data distribution.
@@ 6 Experiments
This section presents evaluations of length deviation using synthetic data to access our theoretical results. An overview of all experiments conducted in this paper can be found in Section 6.1.
Experiment setup. The data generation procedure is described in Appendix C.1. All experiments employ linear models trained with SGD for one epoch using a batch size of $64$. Learning rates are selected via successive halving over the range $10^{-5},1]$. We evaluate miscoverage levels $\alpha\in\{0.01,0.025,0.05,0.075,0.1,0.125,0.15,0.175,0.2\}$. Reported results are averaged over $20$ independent trials, and length deviations are computed on $2000$ test samples.
@@
    Together with (23), the estimated coefficient $b_{1}=-2.24$ (Fig. 3d) implies that $\Delta\sim\alpha^{-2.24}$. This aligns with the theoretical upper bound of order $\mathcal{O}(\alpha^{-2})$. Appendix C.2 provides an additional verification for the existence of this regime.
@@
6.1 Roadmap of Experiments
We here outline the structure of all experiments conducted in the paper.
Synthetic experiments.
Figure 3 in Section 6 and Figure 5 in Appendix C.2 assess the theoretical results developed in this paper. Appendix C.3 further examines optimization effects: Figure 6 investigates SGD with heavy-ball momentum, and Figure 7 reports the case of AdamW (loshchilov2019decoupled). In Appendix C.4, Figure 8 presents results under nonlinear conditional quantile functions. Finally, in Appendix C.5, Figures 910 evaluate alternative convex loss models.
Real-world experiments.
In Appendix D.2.2, Figure 11 presents an empirical evaluation of length deviation of CMR and CQR under different optimizers on five real-world datasets, comparing SGD, SGD with momentum, Adam, and AdamW. In Appendix D.2.3, Figure 12 evaluates non-linear models. Appendix D.3 empirically investigates data-allocation strategies in Figure 13. Appendix D.4 visualizes the distributional properties of the real-world datasets used in our experiments.
7 Limitations, Discussion, and Future Work
Oracle intervals may not be optimal under certain distributions.
Our theoretical analysis shows that the prediction sets produced by CQR and CMR converge to the oracle quantile interval (16) as the training and calibration sample sizes $n$ and $m$ grow. However, the oracle interval itself is not always efficiency-optimal. It is optimal only when the condition in Remark 3.4 holds, which depends on the structure of the conditional distribution. For instance, when the conditional density is multimodal or basin-shaped, the optimal prediction set is not a single interval. In such cases, the prediction sets produced by standard conformal methods such as CMR and CQR do not approximate the optimal set. This limitation stems inherently from the standard non-conformity scores, which are restricted to producing single-interval prediction sets and therefore cannot capture complex distributional structures. One way to improve efficiency in these settings is to move beyond fixed score functions and consider parameterized nonconformity scores that adapt to the data. For instance, recent work such as braun2025minimum employs an optimization-driven framework targeting volume minimization to learn the parametrization. Such approaches could potentially learn transformations that adapt to complex conditional distributions, leading to more efficient prediction sets. This is a promising direction for future research.
Role and limitations of the linearity assumption.
Our theoretical analysis builds on the linearity assumption of the conditional quantiles. This assumption is standard in the theoretical analysis of quantile regression (koenker2005quantile; pan2021multiplier; shen2024online), as it ensures convexity of the objective and therefore the consistency of the SGD estimator as the training data size $n$ grows. While relaxing this assumption is in principle possible, it typically requires additional assumptions on the complexity of the function class or on the estimation error bounds, which may be difficult to verify in practice (bars2025volume).
@@ 8 Conclusion
Appendix A Proofs of Results in CQR